# Making Interpretable Discoveries from Unstructured Data: A High-Dimensional Multiple Hypothesis Testing Approach

### v1

@@ 4 Empirical Applications
Bursztyn et al. [2023] study how the provision of “social cover” affects willingness to publicly dissent to socially stigmatized causes, and the perception of this dissent. As a key application of their theory, Bursztyn et al. [2023] run an information treatment experiment (Experiment 2) in which participants are told they have been matched with another (fictional) respondent that chose to join a campaign to defund the police (a plausible expression of dissent in liberal American politics at the time the experiment was conducted), and then show the participant a tweet that the matched respondent is said to have agreed to publicly post. This tweet has been randomized to either include social cover or no social cover, namely whether or not the tweet indicated that the matched respondent joined the campaign after reading an evidence-based article in support of it (the no social cover condition) or before (the social cover condition). [^4] See Online Appendix Figure B.3 of Bursztyn et al. [2023] for a schematic of the experimental design.
@@
An important outcome collected by this experiment is the participant’s open-ended text response to the question “Why do you think your matched respondent chose to join the campaign to oppose defunding the police?” This open-ended text response is meant to capture the causal effect of social cover on the perception of dissent. As such, this is a setting where discovery is of interest; ideally, to form a holistic understanding of how social cover affects perception of dissent, we do not want to pre-specify what aspects of perception we are interested in, and want to discover any interpretable systematic differences that exist in the text responses across treatment and control groups. In order to make progress on this analysis of unstructured data, Bursztyn et al. [2023] compute a Pearson’s $\chi^{2}$ statistic for all phrases of up to three words per Gentzkow and Shapiro [2010], which they use as an index to rank the phrases that are most differentially expressed in each condition’s open-ended responses. [^5] Specifically, it is a Pearson’s $\chi^{2}$ statistic for a null hypothesis that the propensity to use a given phrase is equal across conditions, per Gentzkow and Shapiro [2010]. We interpret these statistics as simply forming an index, however, because the results of these hypothesis tests are not reported, and no multiple hypothesis testing corrections are implemented. The interpretability of the results from this analysis, is, naturally, hindered by the coarseness of the featurization of the text as three word phrases, as well as the fact that there are no obvious estimands or inferential guarantees. The only qualitative conclusion gleaned by Bursztyn et al. [2023] from this quantitative exercise is: “we find that respondents in the Cover condition are more likely to use phrases related to the article or the associated evidence—for example, ‘article,’ ‘read,’ ‘convincing,’ or ‘increase in crime.’ ” [^6] The Online Appendix Table B.11 of Bursztyn et al. [2023] contains the top ten characteristic phrases in each condition based on the $\chi^{2}$ index.
@@
Can we make more (and more interpretable) discoveries with the framework of this paper? To investigate, we implement a Dict function based on the Gemma Scope autoencoders for Google’s Gemma 2 2B [Lieberum et al., 2024], and specifically the SAE trained on the residual stream of layer 12.777It is recommended to work with SAEs trained on the middle layer of a LLM to capture coarser concepts and features, which is plausibly more suitable for analyzing broad themes in text. We pull the Gemma Scope project’s autointerpretation descriptions of these features from Neuronpedia, [^8] See https://www.neuronpedia.org/gemma-scope for more. though we also implement a custom, “local” autointerpretation pipeline based on the best practices of Paulo et al. [2024], discussed further in Appendix Section 6.2. These are mainstream choices for a SAE and LLM pairing in machine learning interpretability research.
This Dict function yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$, where $n=1033$ and p=12,005p=12,\!005 (non-degenerate features). The hypotheses of interest are, for all $j\in[p]$,
$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[H_{i}Y_{ij}]:=E[X_{ij}]=0,\quad H_{i}=\frac{D_{i}-\pi}{\pi(1-\pi)}.$
@@
for $\pi=0.5$, and where $\{D_{i}=1\}$ indicates the social cover condition. As a baseline, we will implement a simple one-step version of the statistical procedure in Theorem R, which will attain a lower bound on the power achieved by the step-wise procedure.
@@
We may first consider whether or not, as anticipated, simply controlling the FWER would be too conservative for the purposes of discovery in this setting. To explore this, we set $k=1$ and $\alpha=0.05$ in the procedure based on Theorem R (equivalently, we implement the high-dimensional FWER controlling procedure of Belloni et al. [2018]), and we obtain only two significant discoveries: a 13.2 percentage point causal effect on the presence of dictionary feature 3518 (DF-3518) and a 13.4 percentage point causal effect on the presence of dictionary feature 3426 (DF-3426). The Gemma Scope autointerprations of these features paint a picture that corroborates the original findings of Bursztyn et al. [2023]: DF-3426 is described as activating on “references to articles” and DF-3426 on “references to news articles and reports.” However, we only discover two features out of over 12,00012,\!000, and not much new insight relative to the original analysis—even if the interpretation of the features discovered was transparent and automatic, by construction.
In Table 1, we make discoveries based on the procedure of Theorem R now setting $k=5$, still with $\alpha=0.05$, and this time using local autointerpretation descriptions per the method Appendix Section 6.2.
Table 1: Discoveries for Experiment 2 of Bursztyn et al. (2023), 5-FWER Control at $\alpha=5\%$
Feature $\widehat{\text{ATE}}$ Description (“Feature activates highly on…”) t-stat
306 0.099 source that claims police defunding increases crime 4.103
1392 0.118 partial or incomplete reading of an article 3.910
1622 0.072 mentions of a prestigious author or institution (e.g., “Princeton professor”) 4.045
1992 0.095 mentions of an article 4.286
3298 0.126 article about defunding the police and its impact on violent crime 4.223
3426 0.134 mentions of an “article” or article reference 4.803
3518 0.132 articles linking police defunding to increased violent crime 5.020
4320 0.081 article or report content 3.999
8486 0.085 defunding‑the‑police political discourse 3.918
8902 0.079 reading an article that influences a person’s beliefs or viewpoints 4.198
12287 -0.130 modal verbs expressing uncertainty or possibility 3.890
Note: Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025]. For bootstrapping, 10,00010,\!000 iterations are performed.
@@
As can be seen, even with control of $k$-FWER at $k=5$, we obtain 11 rejections of the null, or 11 discoveries. By guarantee of Theorem R, the probability that 5 or more of these 11 discoveries is false is less than 5%. The previous features discovered with the FWER controlling procedure appear, as well as others that represent concepts both absent from the FWER controlled analysis and the original analysis in Bursztyn et al. [2023]. Though there are features that appear to activate on redundant concepts, in totality these discoveries contribute to entirely new characterizations of the causal effect of social cover on dissent. For example, DF-1622 responds to “mentions of a prestigious author or institution” and increases by 7 percentage points in the treatment group, indicating that, as Bursztyn et al. [2023] posit, the social cover mechanism plausibly relies on the perceived credibility of the article cited. DF-1392 increases by 12 percentage points in the treatment group, and responds to discussion of “partial or incomplete reading of an article,” suggesting that that social cover in this experiment may have a negatively valenced effect related to perceptions of ignorance. DF-12287 responds to “modal verbs expressing uncertainty or possibility,” and its probability of activation is decreased by 13 percentage points in the treatment group. This supports the idea that the causal effect of not having social cover is, to first order, generating uncertainty and speculation about dissent, as opposed to an alternative where participants coalesce around a single narrative in the absence of cover. This insight is also congruous with the observation that few large negatively signed causal effects are discovered.
@@
Importantly, all of this analysis was performed automatically, in a handful of minutes on a single A100 GPU in Google Colab. Though this analysis was not preregistered, it is possible to cheaply and automatically replicate these results, and assess their sensitivity to different autointerpretation strategies, dictionary learning methods, and LLMs. Notably, the space of possible concepts was not pruned in any way prior to this analysis, nor were they weighted in some way towards topics related to articles; the analysis was automatic and without human intervention, with no room for motivated data mining.
@@ 4.2 Stantcheva [2024]
Stantcheva [2024] conducts surveys on representative samples of the United States population investigating attitudes towards inflation. These surveys include open-ended text responses to a variety of questions, with the goal of discovering attitudes and opinions that the researcher should not (for concern about priming respondents) or could not (for lack of imagination) pre-specify. [^9] See, e.g., Haaland et al. [2024] for more discussion of the benefits of open-ended survey questions for understanding economic behavior.
@@
An important open-ended prompt that Stantcheva [2024] solicit an answer to is “High inflation is caused by…” In the original analysis, Stantcheva [2024] code open-ended responses with a keyword-based method described in Ferrario and Stantcheva [2022], in which the researcher creates a list of topics and associated keywords in a discretionary way, ranging from “manual to semi-supervised or unsupervised” means. The result of applying this analysis to the open-ended responses to the above prompt yields Figure 3 in Stantcheva [2024], which finds, among other things, that mentions of “Biden and the administration,” “Greed,” “Monetary policy,” “Fiscal policy,” “War and foreign policy,” “Demand vs supply,” and “Supply-side mechanisms (other than input prices)” appear in more than 5% of all responses.
@@
Do we make the same discoveries automatically if we apply the proposed framework instead? Do we make more and new interpretable discoveries? To find out, we use the same models as in Section 4.1, but reduce the space of features by half, filtering out all features that, in the corpus on which the Gemma Scope SAEs were trained, had greater than median empirical token activation frequency. This choice of dimensionality reduction is meant to screen out features that activate in many texts across domains, e.g., features related to grammatical aspects of text, which was handled implicitly in the previous causal analysis by virtue of differencing. (Ideally, such a dimensionality reduction choice would be preregistered, to prevent cherry-picking by filtering.) This yields a dataset $\{X_{ij}\}$ with $n=503$ and $p=3915$ non-degenerate features—still a very high-dimensional inference problem.
We test the $p=3915$ hypotheses $H_{0,j}:E[X_{ij}]=E[Y_{ij}]=0$, i.e., we test the probability of a given feature activating in the population of open-ended responses to the above prompt in the United States. Specifically, by again using 5-FWER control with $\alpha=0.05$ based on the procedure of Theorem R, we make 818 interpretable discoveries. In order to better present these discoveries, using test inversion we produce simultaneous “generalized” confidence sets [Romano and Wolf, 2007], which asymptotically guarantee that the probability that $k$ or more estimands do not lie in the set is less than $\alpha$.
@@
In Table 2, we report the simultaneous confidence intervals for all discoveries with a lower confidence bound above 20%. Notably, we discover, automatically, with no manual pruning or interpretation required, many sensible commonplace themes related to the topic at hand: DF-9804 responds to economic terms, DF-14304 responds to economic terminology, and DF-13447 responds to mentions of inflation. However, we also recover many of the topics from the original analysis in Stantcheva [2024] without any human discretion or intervention: DF-4192 responds to discussion of monetary policy, DF-5719 responds to political commentary, and DF-14747 responds to discussion of supply-side issues. Undoubtedly, many other topics are recovered as we look deeper down the list of discoveries, beyond the top 11 out of 818. However, even in this top 11, we learn that there is a great deal of pure uncertainty expressed in these open-ended responses based on the activation of DF-4794; that macroeconomic indicators get frequently discussed by virtue of DF-11036; and that money supply and government spending and budgeting are often mentioned, per DF-104 and DF-8316, respectively.
Table 2: Largest Discoveries for “High inflation is caused by…” in Stantcheva (2024)
Feature Description (“Feature activates highly on…”) Sim. CI, Lower Sim. CI, Upper
104 excess money supply 0.208 0.352
4192 monetary policy and central‑bank terminology 0.217 0.363
4794 expressions of uncertainty or hedging 0.210 0.355
5719 political commentary about government or the economy 0.843 0.942
8316 government budget or spending 0.279 0.433
9804 economic terms related to the economy 0.268 0.420
11036 macroeconomic indicators 0.212 0.357
13447 inflation or rising prices 0.393 0.553
13574 negative economic conditions 0.230 0.378
14304 economic terminology 0.300 0.456
14747 supply‑related terms 0.234 0.382
Note: Generalized confidence intervals based on $k$-FWER control at $k=5$, $\alpha=5\%$. Descriptions generated using the method described in Appendix Section 6.2, using GPT-OSS 20B [OpenAI, 2025]. For bootstrapping, 10,00010,\!000 iterations are performed.
@@
Once again, this procedure was cheap and fast to implement, and could be replicated and analyzed for sensitivity by any other researcher, quickly. Beyond filtering out greater than median activating dictionary features, no other choices were made to delimit the space of possible discoveries, or focus them in some way on the space of economic- or inflation-relevant concepts; the exact same models that yielded the results of Section 4.1 yielded the results in this section.
@@ 5 Conclusion
Existing literature in empirical economics and econometrics has long suggested the importance of open-ended discovery from high-dimensional or unstructured data. The framework proposed in this paper shows how new statistical procedures for high-dimensional multiple hypothesis testing, when combined with the latest innovations in interpretability methods for machine learning models, can facilitate open-ended, interpretable discovery at scale with higher practicality and higher fidelity than previously possible.
Importantly, the automaticity of the proposed framework makes it subject to very few researcher degrees of freedom, making it resilient to cherry-picking and motivated data-mining without compromising the purpose of discovery in the first place. This is especially true if one couples this framework with minimal preregistration efforts, e.g., simply publicly declaring what dictionary learning model, LLM, and automatic interpretation method are to be used for the analysis of a text, and if any of the features are to be filtered on principled grounds.
A reasonable possible criticism of the proposed framework is that the interpretations associated with dictionary learning methods are not credible. On this topic exactly, there has been an immense amount of attention from AI/ML research communities, which has, to date, primarily concluded that such dictionary features are useful and reliable for discovery, even if they do not uncover “atomic” or fundamental conceptual units in deep neural networks like LLMs [Leask et al., 2025, Peng et al., 2025]. The best practices and methods for moving from features to feature descriptions is an open and important question in the literature, and choices made on this front certainly affect interpretations of discoveries made by the proposed framework. However, as the empirical examples of the previous section make clear, straightforward automatic interpretation methods—which are completely data-driven, replicable, and not subject to motivated human reasoning—yield sensible outputs, scale well with the size of selected sets, can be further scrutinized by humans to assess their validity, and can be perturbed or modified cheaply to assess sensitivity to autointerpretation prompts. Future research related to this framework would seek to apply quantitative methods for validating feature descriptions at scale as well [Paulo et al., 2024, Movva et al., 2025].

### v2

@@ 4 Empirical Applications
Mirroring Example 1, Bursztyn et al. [2023] study how the provision of “social cover” affects willingness to publicly dissent to socially stigmatized causes, and the perception of this dissent. As a key application of their theory, Bursztyn et al. [2023] run an information treatment experiment (Experiment 2) in which participants are told they have been matched with another (fictional) respondent that chose to join a campaign to defund the police (a plausible expression of dissent in liberal American politics at the time the experiment was conducted), and then show the participant a tweet that the matched respondent is said to have agreed to publicly post. This tweet has been randomized to either include social cover or no social cover, namely whether or not the tweet indicated that the matched respondent joined the campaign before reading an evidence-based article in support of it (the no social cover condition) or after (the social cover condition). [^14] See Online Appendix Figure B.3 of Bursztyn et al. [2023] for a schematic of the experimental design.
@@
An important outcome collected by this experiment is the participant’s open-ended text response to the question “Why do you think your matched respondent chose to join the campaign to oppose defunding the police?” This open-ended text response is meant to capture the causal effect of social cover on the perception of dissent. As such, this is a setting where discovery is of interest: ideally, to form a holistic understanding of how social cover affects perception of dissent, we do not want to pre-specify what aspects of perception we are interested in, and want to discover any interpretable systematic differences that exist in the text responses across treatment and control groups. In order to make progress on this analysis of unstructured data, Bursztyn et al. [2023] compute a Pearson’s $\chi^{2}$ statistic for all phrases of up to three words per Gentzkow and Shapiro [2010], which they use as an index to rank the phrases that are most differentially expressed in each condition’s open-ended responses. [^15] Specifically, it is a Pearson’s $\chi^{2}$ statistic for a null hypothesis that the propensity to use a given phrase is equal across conditions, per Gentzkow and Shapiro [2010]. We interpret these statistics as simply forming an index, however, because the results of these hypothesis tests are not reported, and no multiple hypothesis testing corrections are implemented. The interpretability of the results from this analysis, is, naturally, hindered by the coarseness of the featurization of the text as three word phrases, as well as the fact that there are no obvious estimands or inferential guarantees. The only qualitative conclusion gleaned by Bursztyn et al. [2023] from this quantitative exercise is: “we find that respondents in the Cover condition are more likely to use phrases related to the article or the associated evidence—for example, ‘article,’ ‘read,’ ‘convincing,’ or ‘increase in crime.’ ” [^16] The Online Appendix Table B.11 of Bursztyn et al. [2023] contains the top ten characteristic phrases in each condition based on the $\chi^{2}$ index.
@@
Can we make more (and more interpretable) discoveries with the framework proposed in this paper? To investigate, we implement a Dict function based on the Gemma Scope autoencoders for Google’s Gemma 2 2B LLM [Lieberum et al., 2024], and specifically the SAE trained on the residual stream of layer 12.171717It is a well-documented observation that SAEs trained on the middle layer of a LLM seem to capture “coarser” concepts and features (see dicussion in, e.g., Templeton et al. [2024]), which is plausibly more suitable for analyzing broad themes in text. We implement an autointerpretation pipeline as discussed in Section 3.4 using as an explainer LLM the lightweight, open-source reasoning model GPT-OSS 20B [OpenAI, 2025]. A held-out evaluation dataset is created by sample splitting, using 10% of the original dataset in Bursztyn et al. [2023].
The hypotheses of interest are, for all $j\in[p]$,
$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E\left[\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}\right]:=E[X_{ij}]=0,$
@@
for $\pi=0.5$, and where the event $\{W_{i}=1\}$ indicates the social cover condition. This yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$, where $n=930$ and $p=11901$ (non-degenerate features). We implement the simple one-step version of the statistical procedure in Theorem R as per Algorithm 1, which attains a lower bound on the power achieved by the step-wise procedure.
@@
We may first consider whether or not, as anticipated, simply controlling the FWER would be too conservative for the purposes of discovery in this setting. To explore this, we set $k=1$ and $\alpha=0.05$ (equivalently, we implement the high-dimensional FWER controlling procedure of Belloni et al. [2018]), and we obtain only two significant discoveries: a 13.2 percentage point causal effect on the presence of dictionary feature 3518 and a 13.4 percentage point causal effect on the presence of dictionary feature 3426. The autointerprations of these features paint a picture that corroborates the original findings of Bursztyn et al. [2023]: feature 3426 is described as activating on mentions of “defunding the police associated with increased violent crime” and feature 3518 on “mention of an article as a source of evidence.” However, we only discover two features out of approximately $12000$, and not much new insight relative to the original analysis—even if the discovery and interpretation of these features was transparent, inexpensive, and automatic. [^18] The off-the-shelf Gemma Scope autointerpretations echo the same: feature 3426 is described as activating on “references to news articles and reports” and feature 3518 on “references to articles.”
Figure 1: Discoveries for Experiment 2 of Bursztyn et al. (2023)
@@
In Figure 1, we make discoveries now setting $k=5$, still with $\alpha=0.05$. As can be seen, even with control of $k$-FWER at small $k=5$, we obtain 9 rejections of the null, or 9 discoveries. By (large sample) guarantee of Theorem R, the probability that 5 or more of these 9 discoveries is false is less than 5%. The previous features discovered with the FWER controlling procedure appear, as well as others that represent concepts both absent from the FWER controlled analysis and the original analysis in Bursztyn et al. [2023]. Though there are features that appear to activate on redundant concepts, in totality these discoveries contribute to a few entirely new characterizations of the causal effect of social cover on dissent. For example, feature 1622 responds to “Prestigious author source reference” and increases by nearly 8 percentage points in the treatment group, indicating that—exactly as Bursztyn et al. [2023] posit—the social cover mechanism plausibly relies on the perceived credibility of the article cited. Feature 1355 responds to “modal verbs expressing uncertainty or possibility,” and its probability of activation is decreased by nearly 12 percentage points in the treatment group. This supports the idea that the causal effect of not having social cover is, to first order, generating uncertainty and speculation about dissent, as opposed to an alternative scenario where participants coalesce around a single narrative in the absence of cover. This insight is also congruous with the observation that few large negatively signed causal effects are discovered. The A-scores for all discoveries range from 0.98 to 0.66, indicating that some descriptions of discoveries deserve more skepticism than others, and that there is much scope for improving autointerp generation beyond small, open-source reasoning LLMs with the prompting strategy of Appendix Section 6.3.
@@
Importantly, all of this analysis was performed automatically, in approximately one hour, on a single A100 GPU in Google Colab. Though this analysis was not preregistered, it is possible to cheaply and automatically replicate these results, and assess their sensitivity to different autointerpretation strategies, dictionary learning methods, and LLMs. Notably, the space of possible concepts was not pruned in any way prior to this analysis, nor were they weighted in some way towards topics related to articles; the analysis was automatic and without human intervention, with no room for motivated data snooping.
@@ 4.2 Stantcheva [2024]
Mirroring Example 2, Stantcheva [2024] conducts surveys on representative samples of the United States population investigating attitudes towards inflation. These surveys include open-ended text responses to a variety of questions, with the goal of discovering attitudes and opinions that the researcher should not (for concern about priming respondents) or could not (for lack of imagination) pre-specify. [^19] See, e.g., Haaland et al. [2024] for more discussion of the benefits of open-ended survey questions for understanding economic behavior.
@@
An important open-ended prompt that Stantcheva [2024] solicit an answer to is “High inflation is caused by…” In the original analysis, Stantcheva [2024] codes open-ended responses with a keyword-based method introduced in Ferrario and Stantcheva [2022] in which the researcher creates a list of topics and associated keywords in a discretionary way, which may range from “manual to semi-supervised or unsupervised” means. The result of applying this analysis to the open-ended responses to the above prompt yields Figure 3 in Stantcheva [2024], which finds, among other things, that mentions of “Biden and the administration,” “Greed,” “Monetary policy,” “Fiscal policy,” “War and foreign policy,” “Demand vs supply,” and “Supply-side mechanisms (other than input prices)” appear in more than 5% of all responses.
@@
Do we make the same discoveries automatically if we apply the proposed framework instead? Do we make more and new interpretable discoveries? To find out, we use the same Dict function and explainer LLM as in Section 4.1, but reduce the space of features by half, filtering out all features that, in the corpus on which the Gemma Scope SAEs were trained, had greater than median empirical token activation frequency. This choice of dimensionality reduction is meant to screen out features that activate in many texts across domains, e.g., features related to grammatical aspects of text, which was handled implicitly in the previous causal analysis by virtue of differencing. (Ideally, such a dimensionality reduction choice would be preregistered, to prevent cherry-picking by filtering.) Again we split the sample randomly to form a held-out evaluation dataset consisting of 10% of the original dataset analyzed in Stantcheva [2024].
We are interested in testing hypotheses $H_{0,j}:E[X_{ij}]:=E[Y_{ij}]=0$, i.e., we test the probability of a given feature activating in the population of open-ended responses to the above prompt in the United States. This yields a dataset $\{X_{ij}\}_{i\in[n],j\in[p]}$ with $n=453$ and $p=3826$ non-degenerate features—still a very high-dimensional inference problem. We now apply 5-FWER control with $\alpha=0.05$ based on the result in Theorem R per Algorithm 1, and we make 744 interpretable discoveries.
@@
In Figure 2, we report point estimates and simultaneous confidence intervals for all discoveries with a generalized simultaneous lower confidence bound above 20%. Notably, we discover, automatically, with no manual pruning or interpretation required, many sensible commonplace themes related to the topic at hand: feature 9804 responds to discussion of the economy; feature 13447 responds to mentions of inflation; and feature 13574 activates on broad discussion of negative economic conditions. However, we also recover many of the topics from the original analysis in Stantcheva [2024] without any human discretion or intervention: feature 4192 responds to discussion of monetary policy; feature 14747 responds to discussion of supply-side issues; feature 104 activates on discussion of excess and overabundance; feature 8316 responds to fiscal policy mentions; and feature 5719 responds to economic criticisms leveled at the current government. Undoubtedly, many other topics are recovered as we look deeper down the list of discoveries, beyond these 11 largest. However, even in this top 11, we learn that there is a great deal of pure uncertainty expressed in these open-ended responses based on the activation of feature 4794, and also that macroeconomic indicators get frequently discussed by virtue of feature 11036.
Figure 2: Ranked Discoveries for “High inflation is caused by…” in Stantcheva (2024)
@@
Once again, this procedure was cheap and fast to implement, and could be quickly replicated and analyzed for sensitivity by any other researcher. Beyond filtering out greater than median activating dictionary features, no other choices were made to delimit the space of possible discoveries, or focus them in some way on the space of economic- or inflation-relevant concepts; the exact same explainer LLM and SAE that yielded the results of Section 4.1 yielded the results in this section.
@@ 5 Conclusion
Existing literature in empirical economics and econometrics has long suggested the importance of open-ended discovery from high-dimensional or unstructured data. The framework proposed in this paper shows how new statistical procedures for high-dimensional multiple hypothesis testing, when combined with the latest innovations in interpretability methods for machine learning models, can facilitate open-ended, interpretable discovery at scale with higher practicality and higher fidelity than previously possible. It also illustrates how treating (automatic) feature interpretation evaluation as a formal statistical inference problem can yield several scientific benefits in the process of making discoveries.
Importantly, the automaticity of the proposed framework makes it subject to very few researcher degrees of freedom, making it resilient to cherry-picking and motivated data snooping without compromising the purpose of discovery in the first place. This is especially true if one couples this framework with minimal preregistration efforts, e.g., simply publicly declaring what dictionary learning model, LLM, and automatic interpretation pipeline (explainer LLM and prompts) are to be used for the analysis, and if any of the features are to be filtered on principled grounds.